OpenStudy (anonymous):

a golden rectangle has side lengths in the ratio about 1: 1618 if the long side of a golden triangle is 35 cm what is the area?

Directrix (directrix):

Where is the decimal point here supposed to be --> 1618 ?

OpenStudy (anonymous):

1.618

Directrix (directrix):

Okay. So, let's find the length of the shorter side of the golden rectangle. They will be similar polygons so, 1/1.618 = x/35 where x is the shorter side

OpenStudy (anonymous):

okay, I got you

Directrix (directrix):

Cross multiply to get: 1.618*x = 1*35 x = ? @shoortaayy Your task is to solve for x. Post what you get, okay?

Directrix (directrix):

Divide 35 by 1.618

OpenStudy (anonymous):

i got 21.6

Directrix (directrix):

I had x = 21.6316. Look at the instructions. Are you told to approximate the answer to the nearest hundredth or nearest integer or anything?

OpenStudy (anonymous):

the nearest tenth , i got 21.6316 and a whole lot of other numbers too

Directrix (directrix):

The area to the nearest tenth OR the length of the shorter side to the nearest tenth? -------------- Area = 35 * 21.63 Area = ? @shoortaayy

OpenStudy (anonymous):

757.05 , rounded would be 757.1

Directrix (directrix):

Yes, 757.1 square cm

OpenStudy (anonymous):

i have another problem @Directrix

Directrix (directrix):

Close this thread and post the new question in a new thread.